# Revisiting the question of what instantaneous normal modes tell us about liquid dynamics

### v1

@@ V Conclusions
  • The linear scaling of the low-frequency INM DOS has been confirmed for all six liquids. This validates the hypothesis that, as far as the fraction of unstable mode is finite, the Debye scaling does not emerge in the INM spectrum, not even in the finite temperature solid phase. Moreover, this linear scaling has been proven to coexist with a cusp-like singularity in the eigenvalue spectrum, showing that, contrary to recent claims in the literature Schirmacher et al. (2022), the two features are not incompatible.
  • The slope of the INM spectrum in the linear scaling regime $a(T)$ grows with temperature in simple liquids, following an Arrhenius-like exponential form, as previously demonstrated in Jin et al. (2024). On the contrary, metallic liquids display an approximate constant slope, that results (at least in the range of $T$ investigated) independent of temperature.
  • Both the fraction of unstable modes, but also the slope of the INM spectrum in the linear scaling regime $a(T)$, display a sharp jump at the first-order melting transition but no signature at the Frenkel dynamical crossover.
@@
  • Against common expectations, crystals at finite temperature display a rather large number of unstable modes, that becomes negligible only at very low temperature. As a consequence, the scaling of the INM spectrum remains linear. At low temperature, a crossover scaling between a solid-like Debye scaling and a linear regime emerges in the INM spectrum, where the crossover frequency scale decreases with $T$.
  • In all liquids considered, the eigenvalue distribution displays a cusp-like singularity at $\lambda=0$. This feature can be explained both using heterogeneous-elasticity theory Schirmacher et al. (2022); Mossa et al. (2023), but also using the soft potential model of liquids Zürcher and Keyes (1997); Xu et al. (2023). Moreover, the temperature dependence of the intermediate slopes and their difference as a function of temperature do not show the behavior previously reported in the literature for a simulated soft sphere liquid model Schirmacher et al. (2022); Mossa et al. (2023).
@@ Acknowledgements.
Appendix 1 Additional data for Argon
In the main text, we have shown the data for Argon only along the constant pressure line with $P=40$ bar. In Fig. A1, we report additional data at other pressures that confirm the findings discussed in the main text.
Figure A1: Panels (a) and (b) show the INM DOS $\rho(\omega)$ and the density of state function $S(\omega)$ for Argon at $1000$ bar as a function of temperature. Panels (c) and (d) show the same quantities at fixed temperature, $T=120$ K, as a function of the pressure.
Appendix 2 Xenon, Nitrogen and more liquid systems
Argon has been taken in the main text as a benchmark example for our analysis. Here, we report additional data for the other liquids discussed in this manuscript. Fig. A2 reports additional data for the INM DOS at different temperatures for Xe, Pb, N2 and CS2. Fig. A3 shows the density of state function at different temperature for Xe and Pb.
Figure A2: INM DOS $\rho(\omega)$ as a function of temperature for Xe (a), Pb (b), N2 (c) and CS2 (d).
Figure A3: Density of state function $S(\omega)$ for Xe (a) and Pb (b) at different temperatures.
Appendix 3 Self-diffusion constant
In Fig. A4 we show the Arrhenius plot for the self-diffusion constant of the six different liquid systems studied in our work. The corresponding activation energies are reported Eq. (19) in the main text.
Figure A4: Arrhenius plot for the self-diffusion constant $D$ as a function of temperature. Different panels correspond to the six different liquids considered in this work. The red dashed lines indicate the Arrhenius fit.
Appendix 4 Slopes of the eigenvalue distribution $\nu(\lambda)$
In Fig. A5, we show the temperature behavior of the fitted stable and unstable slopes discussed in the main text for the other systems, including Ar, Xe, Pb and Ga.
Figure A5: The “slopes” of the eigenvalue distribution $\upsilon(\lambda)$ as a function of temperature for Ar (a), Xe (b), Pb (c) and Ga (d).

### v2

@@ V Conclusions
  • The linear scaling of the low-frequency INM DOS has been confirmed for all six liquids. This validates the hypothesis that, as far as the fraction of unstable mode is finite, the Debye scaling does not emerge in the INM spectrum, not even in the finite temperature solid phase. This is consistent with the idea that the linear scaling arises from overdamped unstable excitations that do not have propagating wave-like nature. Moreover, this linear scaling has been proven to coexist with a cusp-like singularity in the eigenvalue spectrum, showing that, contrary to recent claims in the literature [36], the two features are not incompatible.
  • The slope of the INM spectrum in the linear scaling regime $a(T)$ grows with temperature in simple liquids, following an Arrhenius-like exponential form, as previously demonstrated in [53]. On the contrary, metallic liquids display an approximate constant slope, that results (at least in the range of $T$ investigated) independent of temperature.
  • Both the fraction of unstable modes, but also the slope of the INM spectrum in the linear scaling regime $a(T)$, display a sharp jump at the first-order melting transition.
@@
  • Against common expectations, crystals at finite temperature display a rather large number of unstable modes, that becomes negligible only at very low temperature. As a consequence, the scaling of the INM spectrum remains linear. At low temperature, a crossover scaling between a solid-like Debye scaling and a linear regime emerges in the INM spectrum, where the crossover frequency scale decreases with $T$. This is caused by the presence of non-diffusive unstable modes that survive in the solid phase. These modes are possibly a spurious effect of the naive INM computation and can be removed using several methods (e.g., [42]).
  • In all liquids considered, the eigenvalue distribution displays a cusp-like singularity at $\lambda=0$. This feature can be explained both using heterogeneous-elasticity theory [36, 45], but also using the soft potential model of liquids [30, 31]. Moreover, the temperature dependence of the intermediate slopes and their difference as a function of temperature do not show the behavior previously reported in the literature for a simulated soft sphere liquid model [36, 45].
  • As discussed in the main text, the linear in frequency DOS obtained from the INM analysis is compatible with the results from $k$-gap theory, suggesting an interesting connection between microscopic normal modes and collective excitations in liquids. Nevertheless, the INM analysis seems to underestimate the Debye quadratic contribution, specially close to the melting temperature. This could be an artifact of the INM approximation and in particular a result of the presence of unstable modes that do not vanish below the melting temperature.